# Not Just Change the Labels, Learn the Features: Watermarking Deep Neural Networks with Multi-View Data

### v1

@@ 6.1 Experimental Settings
Datasets. Building upon prior studies [2, 20], we employ the CIFAR10 [23] and CIFAR100 [23] datasets. Furthermore, we extend our evaluation to include ImageNet [8], a large-scale image dataset which poses challenges to current watermarking techniques. Following [2], we employ half of the training dataset as the source data, while the remaining half serves as the surrogate dataset.
Baselines. We compare MAT with the following baseline approaches,
@@
Metrics. We evaluate the accuracy of the surrogate model $\hat{M}_{\hat{\theta}}$ on trigger set $D_{t}$, and report the p-value using a two-sample t-test, where a small p-value indicates distinguishability between stolen and benign models. Additionally, we report the clean accuracy of both the source and surrogate models.
Attack methods. We consider several commonly used attack methods,
  • Soft-label model extraction attack. The attacker leverages the predictions of the source model on the surrogate model as in Equation 2.
  • Hard-label model extraction. Instead of depending on the soft labels $M_{\theta}(x)=(M^{1}_{\theta}(x),M^{2}_{\theta}(x),...,M^{K}_{\theta}(x))$, the attacker exploits the hard label $y_{hard}=\operatorname*{arg\,max}_{k}{M^{c}_{\theta}(x)}_{k=1}^{K}$. The hard label will be converted into a one-hot representation, which then replaces the soft label in Equation 2 for model extraction attack.
@@
  • Fine-pruning [26]. Given the access to the source model, the attacker prunes the neurons on the last convolutional layer of the source model through the activation level on a small batch of clean data, and then performs fine-tuning.
Implementation details. We utilize ResNet18 [16] as the source and surrogate models for CIFAR10 and CIFAR100, and ViT-base-patch16-384 [11] for ImageNet. Additional results with alternative architectures are provided in the Supplementary due to space limitations. We assume that all surrogate models are randomly initialized, as attackers do not have access to the parameters of the source model. In all the experiments, the size of the trigger set is 100. We first train a clean model on the source data for 200 epochs for extracting multi-view data for constructing the trigger set. Then a source model is trained on the clean data and the trigger set data as in Equation 9. We follow the same training strategies as in [20]. The source model is trained for 200 epochs with stochastic gradient descent with an initial learning rate of 0.1. A linear decay is used to decay the learning rate every 50 epochs. For the details of various attack method, please see the Supplementary. The code is attached in the supplementary materials.
@@ 6.2 Results on Functionality Stealing
Tables 1, 2 and 3 present the source accuracy (Source Acc.), surrogate accuracy (Surro. Acc.), trigger set accuracy (Trig. Acc) and p-value for all the methods on CIFAR10, CIFAR100 and ImageNet, respectively. For the statistical testing based methods - DI and EEF, we only report the p-value. It is evident that following a model extraction attack, all the baselines exhibit a substantial number of errors on the trigger set, posing a significant challenge to ownership verification. Specifically, in the case of a hard-label model extraction attack, none of the baselines can transfer the prediction of the source model on the trigger set to the surrogate model, rendering ownership verification impossible.
@@
Table 1: Our proposed MAT achieves the best watermarking performance on CIFAR10 compared with the baselines. Even after hard-label model extraction attack, MAT still achieves 56% accuracy on the trigger set.
@@
Table 2: Our proposed MAT achieves the best watermarking performance on CIFAR100 compared with the baselines after soft-label model extraction attack. With the hard-label model extraction attack, the attack itself becomes more challenging, yet MAT still achieves higher trigger accuracy compared to other methods.
In contrast, the proposed MAT effectively retains the majority of predictions on the trigger set. Specifically, when compared to the margin-based method, the trigger set accuracy experiences a substantial improvement, rising from 46% to 74%. Even when subjected to a hard-label model extraction attack, our MAT still maintains a trigger set accuracy of 56% on CIFAR10. On CIFAR100, all baseline methods fail to maintaining trigger set accuracy following a soft-label model extraction attack, whereas our proposed MAT demonstrates an accuracy of 77%. Similarly, MAT achieves much higher trigger set accuracy and lower p-value on ImageNet which shows that it can scale to large image datasets.
Consistent with findings in [2], we note that all methods fall short in preserving trigger set accuracy post hard-label model extraction attack. This phenomenon may be attributed to the inherent difficulty of CIFAR100 dataset and ImageNet and the low accuracy of the source model on the surrogate dataset. We also show the results of MAT (no reg.) which does not use feature regularization during training. MAT (no reg.) achieves higher clean accuracy but lower trigger set accuracy compared with MAT. Thus, MAT (no reg.) can be employed when achieving high clean accuracy is more important than ownership verification.
@@ 6.3 Results with White-box Attacks
Model extraction attack is a black-box attack which assumes that the attacker cannot have access to the parameters of the source model. In some cases, the attacker may even know the parameters of the source model in which case the attacker can perform white-box attack. In this section, we examine two white-box attacks, namely fine-tuning and fine-pruning. As depicted in Figure 3, our proposed MAT demonstrates the ability to sustain a reasonably high trigger set accuracy, even when subjected to white-box attacks. This represents a significant strength of MAT in contrast to margin-based approaches, which lack the capability to effectively defend against such attacks.
@@
Figure 3: MAT can achieve much better watermarking performance compared with the margin-based approach with white-box attacks.
@@ 6.4 Ablation Studies
Heterogeneous Surrogate Dataset & Model. MAT does not need to assume that the surrogate data possess the features of the source dataset. In this section, we address a more realistic scenario where the attacker lacks knowledge of the distribution of the source data or the architecture of the source model. Instead of the CIFAR10 dataset, we employ the SVHN dataset [34] as the surrogate dataset, maintaining ResNet18 as the architecture for the surrogate model. Additionally, we explore the use of the CIFAR10 dataset as the surrogate dataset while employing VGG11 [39] as the architecture for the surrogate model. Table 4 demonstrates that in both cases, the proposed MAT consistently outperforms the margin-based approach in terms of trigger set accuracy. Visualization results using T-SNE [31] can be found in the Supplementary.
@@
Different trigger set selection strategies In this section, we further demonstrate the effectiveness of the proposed trigger set selection strategy in comparison to alternative approaches. Specifically, we examine the following strategies:
  • Random selection: This strategy randomly select a subset of samples from the source dataset to form the trigger set.
  • Highest confidence (Highest conf.): This strategy selecgs samples with the smallest logit margin loss over those with the largest.
For both of these strategies, we maintain consistency with the labeling strategy employed by MAT. The results presented in Table 5 and 6 show the performance of these strategies alongside MAT. With the adoption of the proposed trigger set selection strategy, MAT significantly outperforms the other two approaches, emphasizing the effectiveness of choosing multi-view data as the trigger set.
@@
Table 5: Results of trigger set selection strategies on CIFAR10.
@@
Table 6: Results of trigger set selection strategies on CIFAR100
Different trigger set labelling strategies In addition to choosing the trigger set, how to modify the label of the trigger set is also critical for transferring the prediction of the source model on the trigger set to the surrogate model. To further show the effectiveness of our labelling strategy, we examine the following two alternatives:
  • Random labelling: Randomly assign one of the labels, excluding the true label, to the selected sample.
@@
For both of these strategies, we maintain consistency with the trigger set selection strategy employed by MAT. The results presented in Table 7 and 8 showcase the performance of these strategies alongside MAT. With the adoption of the proposed trigger set labelling strategy, MAT significantly outperforms the other two approaches. In particular, MAT assigns the label to the sample in the trigger that the model finds most confusing, making it easier for the model to learn the distinctive features of the assigned classes. Consequently, it becomes feasible to transfer the model’s predictions on the trigger set to the surrogate model.
Clean Acc. Trig. Acc. (%)
@@
Table 7: Results of trigger set labelling strategies on CIFAR10.
Clean Acc. Trig. Acc. (%)
@@
Table 8: Results of trigger set labelling strategies on CIFAR100.
Effect of feature regularization While the inclusion of the feature regularization loss enhances feature learning on the trigger set, it is possible that it may compromise clean accuracy. In Figure 4, we explore the impact of varying the balance parameter $\alpha$ within the range of 0.0, 0.01, 0.05, 0.1. The findings indicate that an increase in $\alpha$ does indeed negatively affect clean accuracy. This can be attributed to the fact that features of different classes are not strictly orthogonal, leading to a decrease in the clean accuracy. Despite this, an increased value of $\alpha$ significantly enhances the effectiveness of the trigger set for watermarking, aligning with our expectations. In practice, to achieve a balance between clean accuracy and watermarking effectiveness, a smaller value of $\alpha$ is preferred to guide the feature learning process. It also worth noting that even without the feature regularization loss, MAT still outperforms the margin-based watermarking method which further demonstrates the effectiveness of the trigger set selection strategy.
[Uncaptioned image] [Uncaptioned image] Figure 4: A large $\alpha$ enhances feature regularization, thereby resulting in improved watermarking performance.
[Uncaptioned image] [Uncaptioned image] Figure 5: MAT can achieve strong watermarking performance with a small trigger set consisting of multi-view data.
Size of the trigger set. In this section, we vary the size of the trigger set, choosing from {50, 100, 150, 200}, in order to assess the impact of trigger size. Figure 5 reveals that varying the trigger set size minimally affects the clean accuracy of both the source model and the surrogate model. However, the accuracy on the trigger set experiences a significant drop as the trigger set size increases. This phenomenon can be easily understood from a feature learning perspective. Due to the limited instances of samples demonstrating multiple features in the datasets, expanding the size of the trigger set marginally integrates additional multi-view data. Taking CIFAR10 as an example, when the trigger size is set to 200, there are 88 correctly classified samples within the trigger set. In comparison, for a trigger set size of 100, there are 75 correctly classified samples. Notably, the absolute number of correctly classified samples remains similar.
@@ 7 Conclusion
In this paper, we propose a novel trigger set-based watermarking method, called MAT, using multi-view data. The proposed MAT is easy to interpret and highly effective for watermarking DNNs. Extensive experiments have been conducted to demonstrate the efficacy of MAT. The findings indicate that MAT not only effectively safeguards against model extraction attacks but also exhibits robustness in the face of whitebox attacks, including fine-tuning and fine-pruning. In summary, MAT stands out as a robust watermarking method, offering enhanced protection for the intellectual property of deep neural networks.

### v2

@@ 6.1 Experimental Settings
Datasets. Building upon prior studies [2, 22], we utilize the CIFAR-10 [25] and CIFAR-100 [25] datasets. Additionally, we extend our evaluation to include ImageNet [8], a large-scale image dataset that challenges existing watermarking techniques. Following [2], we use half of the training dataset as the source data, with the remaining half serving as the surrogate dataset.
Baselines. We compare MAT with the following baseline approaches:
@@
Metrics. We evaluate the accuracy of the surrogate model $\hat{M}_{\hat{\theta}}$ on the trigger set $D_{t}$, and report the p-value using a two-sample t-test, where a small p-value indicates distinguishability between the stolen and benign models. Additionally, we report the clean accuracy of both the source and surrogate models.
Attack methods. We consider several commonly used attack methods:
  • Soft-label model extraction attack. The attacker leverages the predictions of the source model on the surrogate model as in Eq. 2.
  • Hard-label model extraction. Instead of depending on the soft labels $M_{\theta}(x)=(M^{1}_{\theta}(x),M^{2}_{\theta}(x),...,M^{K}_{\theta}(x))$, the attacker exploits the hard label $y_{hard}=\operatorname*{arg\,max}_{k}{M^{c}_{\theta}(x)}_{k=1}^{K}$. The hard label will be converted into a one-hot representation, which then replaces the soft label in Eq. 2 for a model extraction attack.
@@
  • Fine-pruning [30]. Given access to the source model, the attacker prunes the neurons in the last convolutional layer based on their activation levels using a small batch of clean data and then proceeds with fine-tuning.
Implementation details. We utilize ResNet-18 [17] as the source and surrogate models for CIFAR-10 and CIFAR-100, and ViT-base-patch16-384 [11] for ImageNet. Additional results with alternative architectures are provided in the Appendix. We assume that all surrogate models are randomly initialized, as attackers do not have access to the parameters of the source model. In all the experiments, the size of the trigger set is 100. We first train a clean model on the source data for 200 epochs for extracting multi-view data for constructing the trigger set. Then, a source model is trained on the clean data and the trigger set data as in Equation 8. We follow the same training strategies as in [22]. The source model is trained for 200 epochs and optimized using SGD [21] with an initial learning rate of 0.1. A linear decay is used to decay the learning rate every 50 epochs. For the details of various attack methods, please see the Appendix.
@@ 6.2 Results on Functionality Stealing
Tables 1, 2 and 3 present the source accuracy (Source Acc.), surrogate accuracy (Surro. Acc.), trigger set accuracy (Trig. Acc) and p-value for all the methods on CIFAR-10, CIFAR-100, and ImageNet, respectively. For the statistical testing based methods - DI [36] and EEF [27], we report only the p-values. It is evident that following a model extraction attack, all the baselines exhibit a substantial number of errors on the trigger set, posing a significant challenge to ownership verification. Specifically, in the case of a hard-label model extraction attack, none of the baselines can transfer the prediction of the source model on the trigger set to the surrogate model, rendering ownership verification impossible.
Table 1: Our proposed MAT achieves the best watermarking performance on CIFAR-10 compared with the baselines. Even after the hard-label model extraction attack, MAT still achieves 56% accuracy on the trigger set.
@@
Table 2: Our proposed MAT achieves the best watermarking performance on CIFAR-100 compared with the baselines after the soft-label model extraction attack. With the hard-label model extraction attack, the attack itself becomes more challenging, yet MAT still achieves higher trigger accuracy compared to other methods.
@@
In contrast, the proposed MAT effectively retains the majority of predictions on the trigger set. Specifically, when compared to the margin-based method, the trigger set accuracy experiences a substantial improvement, rising from 46% to 74%. Even when subjected to a hard-label model extraction attack, MAT still maintains a trigger set accuracy of 56% on CIFAR-10. On CIFAR-100, all baseline methods fail to maintain trigger set accuracy following a soft-label model extraction attack, whereas our proposed MAT demonstrates an accuracy of 77%. Similarly, MAT achieves much higher trigger set accuracy and a lower p-value on ImageNet, demonstrating its scalability to large image datasets.
Consistent with the findings in [2], we observe that all methods struggle to preserve trigger set accuracy following a hard-label model extraction attack. This phenomenon may be attributed to the inherent challenges of the CIFAR100 dataset and ImageNet, coupled with the low accuracy of the source model on the surrogate dataset. We also present results for MAT (no reg.), which does not use feature regularization during training. MAT (no reg.) achieves higher clean accuracy but lower trigger set accuracy compared to MAT. Thus, MAT (no reg.) can be employed when prioritizing high clean accuracy over ownership verification.
@@ 6.3 Results with White-box Attacks
A model extraction attack is a black-box attack that assumes that the attacker cannot have access to the parameters of the source model. In some cases, the attacker may even know the parameters of the source model in which case the attacker can perform a white-box attack. In this section, we examine two white-box attacks, namely fine-tuning and fine-pruning. As illustrated in Fig. 3, our proposed MAT demonstrates the ability to sustain a reasonably high trigger set accuracy, even when subjected to white-box attacks. This represents a significant strength of MAT in contrast to margin-based approaches, which lack the capability to effectively defend against such attacks.
@@
Figure 3: MAT significantly outperforms the margin-based approach in watermarking effectiveness when facing white-box attacks.
@@ 6.4 Ablation Studies
Heterogeneous Surrogate Dataset & Model. MAT does not need to assume that the surrogate data possess the features of the source dataset. In this section, we address a more realistic scenario where the attacker lacks knowledge of the distribution of the source data or the architecture of the source model. Instead of the CIFAR-10 dataset, we employ the SVHN dataset [38] as the surrogate dataset, maintaining ResNet-18 as the architecture for the surrogate model. Additionally, we explore the use of the CIFAR-10 dataset as the surrogate dataset while employing VGG11 [43] as the architecture for the surrogate model. Table 4 reports that in both cases, the proposed MAT consistently outperforms the margin-based approach in terms of trigger set accuracy. Visualization results using t-SNE [35] can be found in the Appendix.
@@
Different trigger set selection strategies. In this section, we further demonstrate the effectiveness of the proposed trigger set selection strategy in comparison to alternative approaches. Specifically, we examine the following strategies:
  • Random selection (Random): This strategy randomly selects a subset of samples from the source dataset to form the trigger set.
  • Highest confidence (Highest conf.): This strategy selects samples with the smallest logit margin loss over those with the largest.
For both of these strategies, we maintain consistency with the labeling strategy employed by MAT. The results presented in Table 6.4 and 6.4 show the performance of these strategies alongside MAT. With the adoption of the proposed trigger set selection strategy, MAT significantly outperforms the other two approaches, emphasizing the effectiveness of choosing multi-view data as the trigger set.
Table 5: Results of trigger set selection strategies on CIFAR-10.
@@
Table 6: Results of trigger set selection strategies on CIFAR-100
@@
Different trigger set labeling strategies. In addition to choosing the trigger set, how to modify the label of the trigger set is also critical for transferring the prediction of the source model on the trigger set to the surrogate model. To further show the effectiveness of our labeling strategy, we examine the following two alternatives:
  • Random labeling (Random): Randomly assign one of the labels, excluding the true label, to the selected sample.
@@
For both of these strategies, we maintain consistency with the trigger set selection strategy employed by MAT. The results presented in Table 6.4 and 6.4 showcase the performance of these strategies alongside MAT. With the adoption of the proposed trigger set labelling strategy, MAT significantly outperforms the other two approaches. In particular, MAT assigns the label to the sample in the trigger that the model finds most confusing, making it easier for the model to learn the distinctive features of the assigned classes. Consequently, it becomes feasible to transfer the model’s predictions on the trigger set to the surrogate model.
Table 7: Results of trigger set labelling strategies on CIFAR-10.
Clean Acc. (%) Trig. Acc. (%)
@@
Table 8: Results of trigger set labelling strategies on CIFAR-100.
Clean Acc. (%) Trig. Acc. (%)
@@
Effect of feature regularization. While the inclusion of the feature regularization loss enhances feature learning on the trigger set, it may compromise clean accuracy. In Fig. 4, we explore the impact of varying the balance parameter $\alpha$ within the range of {0.0, 0.01, 0.05, 0.1}. The findings indicate that an increase in $\alpha$ does indeed negatively affect clean accuracy. This can be attributed to the fact that features of different classes are not strictly orthogonal, leading to a decrease in clean accuracy. Despite this, an increased value of $\alpha$ significantly enhances the effectiveness of the trigger set for watermarking, aligning with our expectations. In practice, to achieve a balance between clean accuracy and watermarking effectiveness, a smaller value of $\alpha$ is preferred to guide the feature learning process. It is also worth noting that even without the feature regularization loss, MAT still outperforms the margin-based watermarking method which further demonstrates the effectiveness of the trigger set selection strategy.
Refer to caption Refer to caption
Figure 4: A large $\alpha$ enhances feature regularization, thereby resulting in improved watermarking performance.
Refer to caption Refer to caption
Figure 5: MAT can achieve strong watermarking performance with a small trigger set consisting of multi-view data.
Size of the trigger set. In this section, we vary the size of the trigger set, choosing from {50, 100, 150, 200}, to assess the impact of trigger size. Fig. 5 reveals that varying the trigger set size minimally affects the clean accuracy of both the source model and the surrogate model. However, the accuracy of the trigger set experiences a significant drop as the trigger set size increases. This phenomenon can be easily understood from a feature-learning perspective. Due to the limited instances of samples demonstrating multiple features in the datasets, expanding the size of the trigger set marginally integrates additional multi-view data. Taking CIFAR-10 as an example, when the trigger size is set to 200, there are 88 correctly classified samples within the trigger set. In comparison, for a trigger set size of 100, there are 75 correctly classified samples. Notably, the absolute number of correctly classified samples remains similar.
@@ 7 Conclusion
In this paper, we propose a novel trigger set-based watermarking method, called MAT, using multi-view data. The proposed MAT is easy to interpret and highly effective for watermarking deep neural networks (DNNs). Extensive experiments have been conducted to demonstrate the efficacy of MAT. The findings indicate that MAT not only effectively safeguards against model extraction attacks but also exhibits robustness in the face of white-box attacks, including fine-tuning and fine-pruning. In summary, MAT stands out as a robust watermarking method, offering enhanced protection for the intellectual property of DNNs.
Acknowledgements. We would like to thank the anonymous reviewers for their helpful comments. This project was supported by a grant from the University of Texas at Dallas.